1. You can perform a variation of the zinc and sulfur reaction we discussed yesterday, by reacting sulfur with powdered lithium instead, as shown below. If you perform the reaction with 15.0 g of lithium and 15.0 g of sulfur:

16 Li(s)  +   S8(s)  →  8 Li2S(s)

  1. What is the limiting reactant? {/10}
  2. What mass of product can you make? {/4}
  3. What mass of excess reactant will be leftover? {/6}
1
Expert's answer
2021-02-05T04:07:13-0500

15g Li × "\frac{1moleLi}{6.941gLi}" = 2.161 moles Li


15g S8 × "\frac{1moleS_8}{256.52gS_8}" = 0.058 moles S8


1). there are 2.161 moles of Li and 0.058 moles of S8

To use all the Li , how many moles of S8 do we need

2.161 mole Li × "\frac{1moleS_8}{16moleLi}" = 0.135 mole S8

To use all the S8 , how many moles of Li do we need

0.058 moles S8 × "\frac{16moleLi}{1moleS_8}" = 0.928 mole Li

We have enough Li to use all the S8 but we dont have enough S8 to use all the Li .

S8 is the limiting reactant .


2). How many moles of Li2S can we use all the S8

0.058 mole S8 × "\frac{8moleLi_2S}{1moleS_8}" = 0.464 moles

How many grams of Li2S is this

= 0.0464 × 45.95

= 21.3208 g

21.3208 g mass of the product we made


3). Al is the excess reactant

how much are left over

Excess reactant = Total reactant - used reactant

Excess reactant = 2.161 mole - 0.928 mole

= 1.233 moles of Li are left

1.233 mole Li × "\frac{6.941g Li}{1moleLi}" = 8.558 g

8.558 g of Li are left over